OpenStudy (anonymous):

the distance between the center of symmetry of a parallelogram and its longer side is equal to 12 cm. The area of the parallelogram is 720 cm^2, and its perimeter is 140 cm. Determine the length of the longer diagonal of the parallelogram

OpenStudy (anonymous):

I don't know what center of symmetry is :o Can someone please enlighten me?

OpenStudy (anonymous):

, the distance between the center of symmetry of parallelogram and its longer side is equal to 12cm, then this means that the distance between the two longer sides of the parallelogram, is equal to 2 x 12 = 24 cm. Now since the area = 720cm^2, the length of the longer side = 720/24 = 30cm. Since the perimeter = 2x(30 + length of shorter side) = 140, The length of shorter side works out to, (140 - 2x30)/2 = 40cm !! which is LONGER than the longer side !!!

OpenStudy (anonymous):

that should be it

OpenStudy (anonymous):

That doesn't answer the question which is determine the length of the longer diagonal

OpenStudy (anonymous):

ok well then i can't help. SORRY

OpenStudy (anonymous):

My prof. may still ask for it if it could be solve with the given question (which has something wrong in it)

OpenStudy (anonymous):

sorry i tried

OpenStudy (kamille):

Can you make a sketch for us?:)

OpenStudy (tkhunny):



OpenStudy (tkhunny):